OpenStudy (anonymous):

A certain store has a fax machine available for use by its customers. The store charges $1.80 to send the first page and $0.40 for each subsequent page. The total price, P, for the faxing x pages can be modeled by the formula . Determine the number of pages that can be faxed for $5.80. A) 3 pages B) 11 pages C) 15 pages D) 46 pages

OpenStudy (anonymous):

come on mandy, you can do it whats the 1st step

OpenStudy (anonymous):

i have only had 3 hours of sleep I was up all night do the same reveiw test thats 76 questions

OpenStudy (anonymous):

now i have to do it again.

OpenStudy (anonymous):

p=0.40(x-1)+1.80 is the formula that did not show up in the problem

OpenStudy (anonymous):

they charge 1.80 then .40 for each other one so p = 1.80 + ?

OpenStudy (anonymous):

is there a way where i can plug in the answers?

OpenStudy (anonymous):

p = 1.80 +0.4x p = 5.8 5.80 = 1.80 + 0.4 x now?

OpenStudy (anonymous):

5.80-1.80=.4x

OpenStudy (anonymous):

10

OpenStudy (anonymous):

10 pages at the price of 0.4 and how many at 1.80?

OpenStudy (anonymous):

remember 1st page is at 1.80 so 1 + 10 =?

OpenStudy (anonymous):

wait 11 oh

OpenStudy (anonymous):

thank you Im really not this dumb i just haven't had any sleep

OpenStudy (anonymous):

yes, you are smart, but you need to pay attention

OpenStudy (anonymous):

anytime you want help let me know

OpenStudy (anonymous):

thank you